# Mastering Text-to-Image Diffusion: Recaptioning, Planning, and Generating with Multimodal LLMs

### 22 Jan 2024 v1

@@ 1 Introduction
To address these limitations, we introduce a new training-free text-to-image generation/editing framework, namely Recaption, Plan and Generate (RPG), unleashing the impressive multimodal reasoning ability of (multimodal) LLMs to enhance the compositionality and controllability of diffusion models. We propose three core strategies in RPG:
@@ 2.2 Text-to-image Generation
In order to produce meaningful and accurate subregions, we need to carefully specify two components for planning region divisions: (i) region parameters: we define that rows are separated by ”;” and each column is denoted by a series of numbers separated by commas (e.g., ”1,1,1”). To be specific , we first use ”;” to split an image into different rows, then within each rows, we use commas to split a row into different regions. See Fig.5 for better comprehension; (ii) region-wise task specifications to instruct MLLMs: we utilize the CoT reasoning of MLLMs with some designed in-context examples to reason out the plan of region division. We here provide a simplified template of our instructions and in-context examples:
1.Tutorial to learn how to split regions correctly …… 2.Task instructions: Your tasks is to give me spilt ratio when i give you a prompt for multi object, you should do this task step by step: (i) Identify the key phrases in user prompt…… (ii) Recaption all the key phrases with more detailed descriptive prompts…… (iii) Plan and generate complementary region split according to the key phrases… 3. In-context examples User Prompt: A silver drill hair girl in black jacket , wearing blue jeans. Reasoning: Let’s think step by step, firstly we parse the key phrase in the User prompt…… 4.Trigger CoT reasoning ability of MLLMs User Prompt: An old man with his dog is looking at a parrot on the tree. Reasoning: Let’s think step by step……
@@
${\bm{x}}_{t-1}=\text{CRD}({\bm{x}}_{t},y^{\text{\text{base}}},\{\hat{y}^{i}\}_{i=0}^{n},\{R_{i}\}_{i=0}^{n},t,s),$ (4)
@@
Then, in each timestep, we deliver the prompt batch into the denoising network and manipulate the cross-attention layers to generate $n+1$ different latents $\{{\bm{z}}^{i}_{t-1}\}_{i=0}^{n}$ in parallel as in Figure 6. This process can be formulated as:
${\bm{z}}_{t-1}^{i}=\text{Softmax}(\frac{(W_{Q}\cdot\mathcal{\phi}(z_{t}))(W_{K}\cdot\mathcal{\psi}(\hat{y}^{i}))^{T}}{\sqrt{d}})W_{V}\cdot\mathcal{\psi}(\hat{y}^{i}),$ (6)
where image latent ${\bm{z}}_{t}$ is the query and each subprompt $y^{i}$ works as a key and value. $W_{Q},W_{K},W_{V}$ are linear projections and $d$ is the latent projection dimension of the keys and queries. Then, we shall proceed with resizing and concatenating the generated regions $\{{\bm{z}}_{t-1}^{i}\}^{n}_{i=0}$, according to their assigned region numbers (from 00 to $n$) and respective proportions. Here we denoted each resized complementary hidden states as:
${\bm{z}}_{t-1}^{i}(h,w)=\text{Resize}({\bm{z}}_{t-1}^{i}),$ (7)
where $h,w$ are the height and the width of the region. We directly concatenate them along the spatial axes:
@@ 3.1 Text-to-Image Generation
Generalizing to Different Diffusion Backbones
Our RPG framework is of great generalization ability, and can be easily generalized to different diffusion backbones. As presented in Figure 10, with Stable Diffusion v2.1 (Rombach et al., 2022) and recent SOTA model ConPreDiff (Yang et al., 2023b), RPG exhibits different performances and better diffusion backbone can generate the images better aligning with the textual semantics (e.g., ”dwarfs”). We also generalize RPG to ControlNet (Zhang et al., 2023a) for incorporating more conditional modalities. As demonstrated in Figure 3, our RPG can significantly improve the composibility of original ControlNet in both image fidelity and textual semantic alignment.
@@ 5 Related Work
Large Language Models (LLMs) (ChatGPT, 2022; Chung et al., 2022; Zhang et al., 2022; Iyer et al., 2022; Workshop et al., 2022; Muennighoff et al., 2022; Zeng et al., 2022; Taylor et al., 2022; Chowdhery et al., 2023; Chen et al., 2023b; Zhu et al., 2023; Touvron et al., 2023; Yang et al., 2023a; Li et al., 2023a) have profoundly impacted the AI community. Leading examples like ChatGPT (ChatGPT, 2022) have showcased the advanced language comprehension and reasoning skills through techniques such as instruction tuning (Ouyang et al., 2022; Li et al., 2023c; Zhang et al., 2023c; Liu et al., 2023). Further, Multimodal Large language Models (MLLMs), (Koh et al., 2023; Yu et al., 2023; Sun et al., 2023b; Dong et al., 2023; Fu et al., 2023; Pan et al., 2023; Wu et al., 2023a; Zou et al., 2023; Yang et al., 2023d; Gupta & Kembhavi, 2023; Surís et al., 2023) integrate LLMs with vision models to extend their impressive abilities from language tasks to vision tasks, including image understanding, reasoning and synthesis. The collaboration between LLMs (ChatGPT, 2022; OpenAI, 2023) and diffusion models (Ramesh et al., 2022; Betker et al., 2023) can significantly improve the text-image alignment as well as the quality of generated images (Yu et al., 2023; Chen et al., 2023b; Dong et al., 2023; Wu et al., 2023b; Feng et al., 2023; Pan et al., 2023). For instance, GILL (Koh et al., 2023) can condition on arbitrarily interleaved image and text inputs to synthesize coherent image outputs, and Emu (Sun et al., 2023b) stands out as a generalist multimodal interface for both image-to-text and text-to-image tasks. Recently, LMD (Lian et al., 2023) utilizes LLMs to enhance the compositional generation of diffusion models by generating images grounded on bounding box layouts from the LLM (Li et al., 2023b). However, existing works mainly incorporate the LLM as a simple plug-in component into diffusion models, or simply take the LLM as a layout generator to control image compositions. In contrast, we utilize MLLMs to plan out image compositions for diffusion models where MLLMs serves as a global task planner in both region-based generation and editing process.

### 06 Feb 2024 v2

@@ 1 Introduction
To address these limitations, we introduce a new training-free text-to-image generation framework, namely Recaption, Plan and Generate (RPG), unleashing the impressive reasoning ability of multimodal LLMs to enhance the compositionality and controllability of diffusion models. We propose three core strategies in RPG:
@@ 2.2 Text-to-image Generation
In order to produce meaningful and accurate subregions, we need to carefully specify two components for planning region divisions: (i) region parameters: we define that rows are separated by ”;” and each column is denoted by a series of numbers separated by commas (e.g., ”1,1,1”). To be specific , we first use ”;” to split an image into different rows, then within each row, we use commas to split a row into different regions, see Figure 5 for better comprehension; (ii) region-wise task specifications to instruct MLLMs: we utilize the CoT reasoning of MLLMs with some designed in-context examples to reason out the plan of region division. We here provide a simplified template of our instructions and in-context examples:
1.Task instruction You are an smart region planner for image. You should use split ratio to specify the split method of the image, and then recaption each subregion prompts with more descriptive prompts while maintaining the original meaning. 2.Multi-modal split tutorial …… 3. In-context examples User Prompt: A girl with white ponytail and black dress are chatting with a blonde curly hair girl in a white dress in a cafe. # Key pharses extraction and Recaption # Split ratio Planning # Composition Logic # Aesthetic Considerations: # Final output 4.Trigger CoT reasoning ability of MLLMs User Prompt: An old man with his dog is looking at a parrot on the tree. Reasoning: Let’s think step by step……
@@
${\bm{x}}_{t-1}=\text{CRD}({\bm{x}}_{t},y^{\text{\text{base}}},\{\hat{y}^{i}\}_{i=0}^{n},\{R^{i}\}_{i=0}^{n},t,s),$ (4)
@@
In each timestep, we deliver the prompt batch into the denoising network and manipulate the cross-attention layers to generate different latents $\{{\bm{z}}^{i}_{t-1}\}_{i=0}^{n}$ and ${\bm{z}}_{t-1}^{\text{base}}$ in parallel, as demonstrated in Figure 6. We formulate this process as:
${\bm{z}}_{t-1}^{i}=\text{Softmax}(\frac{(W_{Q}\cdot\mathcal{\phi}({\bm{z}}_{t}))(W_{K}\cdot\mathcal{\psi}(\hat{y}^{i}))^{T}}{\sqrt{d}})W_{V}\cdot\mathcal{\psi}(\hat{y}^{i}),$ (6)
where image latent ${\bm{z}}_{t}$ is the query and each subprompt $\hat{y}^{i}$ works as a key and value. $W_{Q},W_{K},W_{V}$ are linear projections and $d$ is the latent projection dimension of the keys and queries. Then, we shall proceed with resizing and concatenating the generated latents $\{{\bm{z}}_{t-1}^{i}\}^{n}_{i=0}$, according to their assigned region numbers (from 00 to $n$) and respective proportions. Here we denote each resized latent as:
${\bm{z}}_{t-1}^{i}(h,w)=\text{Resize}({\bm{z}}_{t-1}^{i},R^{i}),$ (7)
where $h,w$ are the height and the width of its assigned region $R^{i}$. We directly concatenate them along the spatial axes:
@@ 3.1 Text-to-Image Generation
Figure 10: Generalizing RPG to different (multimodal) LLM architectures, including Llama 2 (Touvron et al., 2023b), Vicuna (Chiang et al., 2023) and MiniGPT-4 (Zhu et al., 2023).
@@
Generalizing to Various LLMs and Diffusion Backbones
Our RPG framework is of great generalization ability, and can be easily generalized to various (M)LLM architectures (in Figure 10) and diffusion backbones (in Figure 11). We observe that both LLM and diffusion architectures can influence the generation results. We also generalize RPG to ControlNet (Zhang et al., 2023a) for incorporating more conditional modalities. As demonstrated in Figure 3, our RPG can significantly improve the composibility of original ControlNet in both image fidelity and textual semantic alignment.
@@ 5 Related Work
Large Language Models (LLMs) (ChatGPT, 2022; Chung et al., 2022; Zhang et al., 2022; Iyer et al., 2022; Workshop et al., 2022; Muennighoff et al., 2022; Zeng et al., 2022; Taylor et al., 2022; Chowdhery et al., 2023; Chen et al., 2023b; Zhu et al., 2023; Touvron et al., 2023a; Yang et al., 2023a; Li et al., 2023a) have profoundly impacted the AI community. Leading examples like ChatGPT (ChatGPT, 2022) have showcased the advanced language comprehension and reasoning skills through techniques such as instruction tuning (Ouyang et al., 2022; Li et al., 2023c; Zhang et al., 2023c; Liu et al., 2023). Further, Multimodal Large language Models (MLLMs), (Koh et al., 2023; Yu et al., 2023; Sun et al., 2023b; Dong et al., 2023; Fu et al., 2023; Pan et al., 2023; Wu et al., 2023a; Zou et al., 2023; Yang et al., 2023d; Gupta & Kembhavi, 2023; Surís et al., 2023) integrate LLMs with vision models to extend their impressive abilities from language tasks to vision tasks, including image understanding, reasoning and synthesis. The collaboration between LLMs (ChatGPT, 2022; OpenAI, 2023) and diffusion models (Ramesh et al., 2022; Betker et al., 2023) can significantly improve the text-image alignment as well as the quality of generated images (Yu et al., 2023; Chen et al., 2023b; Dong et al., 2023; Wu et al., 2023b; Feng et al., 2023; Pan et al., 2023). For instance, GILL (Koh et al., 2023) can condition on arbitrarily interleaved image and text inputs to synthesize coherent image outputs, and Emu (Sun et al., 2023b) stands out as a generalist multimodal interface for both image-to-text and text-to-image tasks. Recently, LMD (Lian et al., 2023) utilizes LLMs to enhance the compositional generation of diffusion models by generating images grounded on bounding box layouts from the LLM (Li et al., 2023b). However, existing works mainly incorporate the LLM as a simple plug-in component into diffusion models, or simply take the LLM as a layout generator to control image compositions. In contrast, we utilize MLLMs to plan out image compositions for diffusion models where MLLMs serves as a global task planner in both region-based generation and editing process.